Stlyez Corp. has a monthly demand of 2,000 units for a product. The product is used at a constant rate over the 365 days. The annual holding cost for the product is estimated to be $4.00 per unit and the cost of placing each order is $150.00. Current order quantity (lot size) is 1000 units.

What is the total annual inventory cost for Stlyez Corp, if it orders as per EOQ?

Select one:

A) $5,196

B) $5,367

C) $1,500

D) $1,549

E) None of the above
Answer #1

1. B. 5367

ANNUAL DEMAND = 24000
ORDERING COST = 150
HOLDING COST = 4

EOQ = SQRT(2 * D * S / H), WHERE D = DEMAND, S = ORDERING COST AND H = HOLDING COST = SQRT(2 * 24000 * 150 / 4) = 1342

ANNUAL HOLDING COST = AVERAGE INVENTORY * PER UNIT HOLDING COST = (1342 / 2) * 4 = 2684

ANNUAL ORDERING COST = NUMBER OF ORDERS * ORDERING COST = (24000 / 1342) * 150 = 2683

COST OF MANAGING = ANNUAL HOLDING COST + ANNUAL ORDERING COST = 2684 + 2683 = 5367

Do leave a like for the answer if you found it helpful.